OpenStudy (anonymous):

If f(x)=x^8+3x^3+3x find the second derivative f"(x) . I need starting to find just the first derivative.

OpenStudy (ranga):

The derivative of x^n is nx^(n-1). Apply this formula for each term in f(x)=x^8+3x^3+3x and you will get the first derivative. Do it again and you will get your second derivative.

OpenStudy (anonymous):

with the 3x using the nx^(n-1) would it just 8x^7+(3)3x^2+(1)3? or +(1)3x? @ranga

OpenStudy (ranga):

Your first two terms are correct. But what is the derivative of 3x? First let us find the derivative of x using the formula derivative of x^n is nx^(n-1) x = x^1 So derivative of x^1 = (1)x^(1-1) = (1)x^0 = 1 So derivative of 3x = 3

OpenStudy (anonymous):

Makes sense...so it's 8x^6+6x^2+3 correct? And then you use the same formula again so would it be (before multiplying) (7)8x^6+(2)6x+3?

OpenStudy (ranga):

Oh, you are already on to the second derivative. But first simplify the first derivative: f'(x) = 8x^7 + (3)3x^2 + 3 f'(x) = 8x^7 + 9x^2 + 3 f''(x) = ? (Note derivative of a constant is 0).

OpenStudy (anonymous):

So would it be 56x^6+18x?

OpenStudy (ranga):

Good!

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (ranga):

You are welcome. Best wishes.